OpenStudy (aaalive):

What is the difference? 1/8 - 3/10

TheSmartOne (thesmartone):

@AAAlive What would the least common denominator be?

OpenStudy (aaalive):

40

TheSmartOne (thesmartone):

ok, so multiply 1/8 in the numerator and denominator by 5 so we can get a final denominator as 40. And for 3/10, multiply the numerator and denominator by 4 to get the denominator as 40. So, tell me what you get (:

OpenStudy (aaalive):

I know that it will be 5/40 - 12/40 but I am not good with negative numbers, so thats where I got caught up

TheSmartOne (thesmartone):

ok 5 - 12 = ?? That's what we need to find :) Let's factor out -1 so it's easier 5 - 12 = (-1)(-5 + 12) = (-1)(12 - 5) = ?? So what is 12 - 5= ?

OpenStudy (aaalive):

7

TheSmartOne (thesmartone):

and (-1)(7) = ?

OpenStudy (aaalive):

-7

TheSmartOne (thesmartone):

and therefore 5 - 12 = -7

OpenStudy (aaalive):

okay so -7/40?

TheSmartOne (thesmartone):

yup :)

OpenStudy (aaalive):

THANKS DOOD

TheSmartOne (thesmartone):

No problem :)